Hello my name im renny 17 i have two math problems to ask u about! Here arem my problems (rational functions)
2y-1/4y = 4/6 (solving rational func)
t^3 - t^2 - 4 / t^2 + t -2 (simplify rational expressions)
Thankyou so much for looking at these problems!!

Answer by Nate(3500)

2y - 1/4y = 4/6
(8/4)y - (1/4)y = 4/6
(7/4)y = 4/6
y = (4/7)(4/6) = (16/42) = (8/21)
y = (8/21) this function is a horizontal line
------------------------------------------------------------------------------
(t^3 - t^2 - 4) / (t^2 + t -2)
((t-2)(t^2+t+2))/((t+2)(t-1))
cannot be simplified futher
N8